Please answer it all clearly Sochi can read it and understand it .
STION Part A Tangle Company produces fidget toys for autism children that sells for RM40 each. For the upcoming year, management expects fixed costs of a total RM440,000 and variable costs to be RM18 per unit. Required (a) Compute the break-even point in units by using the mathematical equation. (3 marks) (b) Compute the break-even point in Ringgit Malaysia (RM) by using the contribution margin (CM) ratio. (c) Compute the margin of safety percentage assuming actual sales is RM1,000,000 (3 marks) (1 mark) (d) Compute the sales required in Ringgit Malaysia (RM) to earn net income of (3 marks) RM330,000 by using the mathematical equation. Part B Followings are information about Beryls Chocolate: 1. Actual production: 4,000 boxes of chocolates 2. Direct materials: actual used 4,300 pounds of chocolate at RM 15 per pound. Standard price is RM 16 per pound and standard quantity allowed per box of chocolates is 1 pound. 3. Direct labour: actual hours of 6,400 at RM 30.5 per hour. Standard hours allowed per box of chocolate is 1.5 hours. Standard price is RM 30 per hour Required Calculate the followings: i. Material price variance ii. Material quantity variance ii. Labour price variance iv. Labour quantity variance (2.5 marks) (2.5 marks) (2.5 marks) (2.5 marks) ITOTAL 20 MARKSI
Answer #1

PART A:-

(a) Break even point in units = Fixed Cost / (Sales price per unit - Variable cost per unit)

= RM 440,000 / (RM 40 - RM 18)

= 20,000 units

(b) Contribution margin =  Sales price per unit - Variable cost per unit = RM 40 - RM 18 = RM 22

Contribution margin ratio = Contribution margin / Sales price per unit = RM 22 / RM 40 = 55%

Break even point in RM = RM 440,000 / 55% = RM 800,000

(c) Margin of safety percentage = (Actual sales - Break even sales) / Actual sales * 100

= (RM 1,000,000 - RM 800,000) / RM 1,000,000 * 100

= 20%

(d) Contribution margin = Sales - Variable cost

We also know that,  Contribution margin = Fixed cost + Profit

= RM 440,000 + RM 330,000 = RM 770,000

Contribution margin = Contribution margin per unit * Unit sales

Contribution margin per unit * Unit sales =  Fixed cost + Profit

Unit sales = Fixed cost + Profit / Contribution margin per unit

= RM 770,000 / (RM 40 - RM 18) = 35000 units

Required sales in RM = 35000 * RM 40 = RM 1,400,000

PART B:-

(i) Actual quantity = 4300 pounds

Actual price = RM 15 per pound

Actual cost = Actual quantity * Actual price = 4300 * 15 = RM 64,500

Standard price = RM 16

Standard cost = Actual quantity * Standard Price = 4300 * 16 = RM 68,800

Material Price Variance = Actual Cost - Standard Cost

= RM 64,500 - RM 68,800

= RM 4300 Favorable

(ii) Material quantity Variance = Standard Price * ( Standard quantity - Actual quantity)

= RM 16 * ( 4000 - 4300 )

= - RM 4800 Adverse

(iii) Labor Price Variance = ( Standard rate - Actual rate) * Actual hours

= ( 30 - 30.5 ) * 6400

= -RM 3200 Adverse

(iv) Labor quantity Variance = Standard rate * ( Standard hours - Actual hours)

= 30 * [ 6000(4000 * 1.5) - 6400]

= - RM 12,000 Adverse